HUBBLE SPACE TELESCOPE DAILY REPORT #2258 PERIOD COVERED: 0000Z (UTC) 11/24/98 - 0000Z (UTC) 11/25/98 Daily Status Report as of 329/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Twelve Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. As described in HSTAR 6784 and yesterday's report, the acquisition for the second and third iterations of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.3 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed FOC/96 6350 (FOC spectroscopy of BALQSOs) The Faint Object Camera (f/96) was used to make ultraviolet spectrographic images of BALQSO 1011+0906 which is known to show broad emission lines from abundant ions. This will help us improve our knowledge of these BAL class objects. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Eight Sets of STIS/CCD 8057 (STIS CCD Performance Monitor {continued}) The Space Telescope Imaging Spectrograph (CCD) was used to monitor to overall CCD performance. As described in HSTAR 6784 and yesterday's report, the acquisition for the second iteration of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.6 Completed WF/PC-2 7344 (Witnessing the Birth of Radio Galaxies: WF/PC-2 & NICMOS Observations of CSOs) The WF/PC-2 was used to make observational images of the cores of bright, nearby galaxy (Compact Symmetric Object) CSO 4C31.04 in order to try to witness the birth of a radio galaxy. Only HST can image these galaxies at the same size scale of the CSO itself, also allowing us to search for nuclear structures aligned with the radio source in this formative stage. As described in HSTAR 6784 and yesterday's report, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.7 Completed Four Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. As described in HSTAR 6784 and yesterday's report, the acquisition for the first iteration of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the proposal completed with no other reported problems. 1.8 Completed Nine Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. As described in HSTAR 6784 and yesterday's report, the acquisition for the first iteration of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further problems were reported. 1.9 Completed WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxy 3C29.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 6626 (Imaging of Z>1 - Massive Star Forming Galaxies) The WF/PC-2 was used to make observations of the massive star forming galaxy SSA22W which is known to have a redshift of Z>1.0. The central regions of the galactic strips have already been imaged with HST and the current observations are to cover the samples required to finish the set. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Five Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 6834 (Lenses, Mirrors, and HST: IRAS FSC10214+4724 Under a Compound Microscope) The WF/PC-2 was used to make observations of IRAS IRASFSC102. Our existing WF/PC-2 images using filter F814W provides compelling evidence that the this object which is at redshift of 2.3 and is gravitationally lensed. The observations were completed as planned, and no anomalies were reported. 1.13 Completed Four Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. Two TTRs were written: one for six data hits and the other for two data inversions. HSTAR 6786 describes a NICMOS single bit error that occurred at 328/214956Z. The NICMOS internal log was dumped at 328/2220Z. HSTAR 6788 documents three additional NICMOS single bit errors occurring at 328/231256Z, at 329/000155Z, and at 329/003255Z. After the latest occurrence, the ground system limit was raised to 53. 3.0 SIGNIFICANT FORTHCOMING EVENTS: No daily report will be produced for the Thanksgiving holiday or on Friday, November 27, 1998. The next daily report will be submitted on Monday, November 30, 1998. Continuation of normal science observations and calibrations. /DMH